OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 8 when x = 6, write the direct linear variation equation.

OpenStudy (anonymous):

f(x)=8+6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 48 when x = 8, find the value of f(x) when x = 2 how about this

OpenStudy (anonymous):

42 because 40 +8 =48 or f(x)=40+8 so f(x)=40+2 or 42

OpenStudy (anonymous):

thats not a option its 12 6 8 3

OpenStudy (anonymous):

sorry its 12 8 times 6 is 48 and 6 times by 2 is 12

OpenStudy (anonymous):

so its 12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which does NOT show a direct variation between x and y? y = 2x y = 0.5x

OpenStudy (anonymous):

or y= y/9 , y=x/9

OpenStudy (anonymous):

y=y/9 because x is nonexistent

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

you are welcome!!

OpenStudy (anonymous):

can you do this one please What is the domain of the function: {(1, 2); (2, 4); (3, 6); (4, 8)}

OpenStudy (anonymous):

that one is one im not so sure ill be able to help you with! i learned that so long ago!